# Li and Zhu

### v1

@@ 2.1 Problem Setup
(a) Desirability
(b) $\epsilon$-optimal uniform policy
(c) $\epsilon$ vs regret
2.2 A Tight Reformulation with Gaussian Processes
To solve Problem $\mathcal{P}^{*}$, it is necessary to impose certain assumptions on the structure of the stochastic process $U(A_{i})$. Specifically, we assume that this process can be modeled as a stationary Gaussian process (GP) with a mean of zero and a covariance function $k(a,a^{\prime})$: {assumption} The qualitative desirability $U(a)$ can be modeled by a stationary GP with a mean of zero and a covariance function $k(a,a^{\prime})$. This assumption is quite general as GPs are widely recognized in the literature for their flexibility and capacity to model a broad range of stochastic processes. By assuming stationarity, we are utilizing the principle of indifference to concretize the notion that since qualitative desirability is unknown, it therefore should not apriori depend on what the action is.
Utilizing Assumption 2.2, we can bound the objective in Problem $\mathcal{P}^{*}$ as followed:
@@ 2.2 A Tight Reformulation with Gaussian Processes
Problem $\mathcal{P}^{*}$ admits the following upper and lower bounds:
@@
The proof is detailed in Appendix 6. By assuming that the qualitative desirability follows a general GP, we derive both lower and upper bounds for the original objective. In the case where $m=1$, the term $E_{m}$ becomes zero, reducing both bounds to the classical optimal policy problem that disregards diversity:
$\max_{\pi}\mathbb{E}_{\pi}[Y(A_{i})].$
This outcome confirms that when the goal is to recommend only a single solution to the decision-maker, the unobservable qualitative desirability can be effectively ignored.
However, when multiple solutions are to be recommended, the objective naturally evolves into a trade-off between quantitative optimality and a second term representing solution diversity. Specifically, $\rho[\pi]$ captures the expected correlation between the qualitative desirability components $U(A_{i})$ and $U(A_{j})$, serving as a measure of similarity or concentration among the generated solutions. When $\rho[\pi]=1$, the solutions are perfectly correlated in terms of their qualitative desirability, indicating a lack of diversity. The term $\sigma\sqrt{1-\rho[\pi]}$ in both bounds therefore incentivizes minimizing this correlation to promote diversity among the solutions.
@@ 2.3 Theoretical Results
    Assume the space of actions $\mathcal{A}$ is $[-q,q]$. If $U(a)$ follows the Gaussian kernel $k(a,a^{\prime})=\exp\left(-\frac{(a-a^{\prime})^{2}}{2h^{2}}\right)$, let:
@@
[…]
@@ 2.4 Incorporating Human Preference Feedback
The framework presented above focuses on a common scenario where the algorithm generates a range of desirable actions for the human decision-maker to select from. We can further extend this framework to incorporate the feedback from human decision-makers, which allows the algorithm to refine its understanding of implicit preferences and enhance the accuracy of future recommendations in a multi-stage process.
@@ 3.1 Deep Generative Approach
Our proposed framework allows a wide range of generative models to directly parameterize the policy $\pi$, enabling their use in solving the optimization problem presented in Problem $\mathcal{P}^{*}$. By harnessing the capabilities of cutting-edge generative models, our approach can produce a rich distribution of actions that is difficult to capture with traditional parametric models. These generative models not only optimize the objective function but also maintain a desirable level of diversity, which is critical in many real-world applications. In the experiments, we focus on a simple yet flexible generative architecture using the reparameterization trick (Kingma 2013) for clarity and ease of implementation. Specifically, instead of sampling $a$ directly from the distribution $\pi$, we express $a$ as a deterministic function of the model parameter $\theta$ and an auxiliary noise variable $\epsilon_{\text{NN}}$ drawn from a standard normal distribution. Thus, the generative process can be written as:
@@ 3.2 Iterative Curation Approach
Despite the success of generative models in various fields, applying these techniques to integer optimization problems – such as routing, assignment, and scheduling – presents unique challenges. These problems are often characterized by discrete and combinatorical structures, where the solution space is non-continuous and highly constrained. Traditional methods in integer programming focus on identifying a single optimal solution, often using techniques such as branch and bound (Land and Doig 1960) or cutting planes (Gomory 1958). These methods are not inherently designed to generate a distribution of solutions, making the direct application of generative curation frameworks less straightforward.
To address this limitation, we propose an iterative approach that enables the generation of diverse solutions over multiple iterations. As illustrated by Figure 4(c), this approach allows the model to incrementally refine the solution set, ultimately converging to a distribution that can approximate the optimal solution distribution for Problem $\mathcal{P}^{*}$. Through each iteration, the generated solutions increasingly reflect the desired balance between diversity and optimality, leveraging the strengths of generative methods while addressing the unique constraints posed by complex optimization problems. Specifically, we consider an iterative approach as outlined in Algorithm 2. At every step, we optimize the sample estimate of the objective function in Equation (3). As shown in Figure 4(c), with well-chosen noise variance $\epsilon^{2}_{\text{DIS}}$ (similar to the approach used for selecting $\epsilon^{2}_{\text{NN}}$), this procedure can empirically converge to the same distribution as our optimal solution to the original problem.
@@ 4 Experiments
To ensure a fair comparison, we rerun each baseline $m$ times and select the action that gives the highest human desirability.
To measure the performance of the algorithms, we adopt total regret as the key evaluation metric in our synthetic experiments, defined as:
@@
To evaluate the robustness of the methods, we also examine the $95\%$ quantile upper bound (Up) and $5\%$ quantile lower bound (Low) of the regret. Furthermore, we measure empirical diversity using $\sqrt{1-\widehat{\rho}[\pi]}$, where $n$ random actions $\{a_{i}\}_{i\leq n}$ are generated from policy $\pi$, and $\widehat{\rho}[\pi]$ is computed as $1/\sigma^{2}\sum_{i=1}^{\lfloor{n/2}\rfloor}k(a_{2i},a_{2i-1})/\lfloor{n/2}\rfloor$.
@@ 4.1 Synthetic Results
Methods $1$D Gaussian $2$D Ackley Knapsack
$\mathbb{E}R(\pi)$ Low Up $\sqrt{1-\widehat{\rho}[\pi]}$ $\mathbb{E}R(\pi)$ Low Up $\sqrt{1-\widehat{\rho}[\pi]}$ $\mathbb{E}R(\pi)$ Low Up $\sqrt{1-\widehat{\rho}[\pi]}$
@@
An example of the synthetic $2$D desirability function in a bounded action space. The transparent surface in red represents the true desirability $\ell$ and the surface in gray represents the quantitative desirability $Y$, which is an $2$D negative Ackley function. In this example, we set $\sigma=10$.
@@
    2D Ackley: We select a negative Ackley function (Ackley 2012) in a bounded $2$D space $[-3,3]\times[-3,3]$ as the quantitative desirability $Y$ and utilize zero-mean exponential-kernel GP with $h=0.5$ for the qualitative desirability. Figure 4.1 illustrates an example of potential desirability functions with $\sigma=10$, where the quantitatively optimal action is $a^{*}=(0,0)$.
@@
    Knapsack: We apply the proposed framework to a discrete setting by solving the classic knapsack problem (Kellerer et al. 2004). This problem consists of a set of $d=10$ items, each characterized by a specific weight and value (randomly generated integers between 00 and $10$), and a knapsack with a defined capacity $20$. The objective is to determine the optimal combination of items that maximizes the total value without exceeding the knapsack’s capacity. We set $\sigma=10$ and model the qualitative desirability using a GP, where the Hamming distance is used to measure the “similarity” between two actions, and an exponential covariance kernel with $h=0.5$ is adopted. Each possible action is represented as a binary vector, where each element indicates whether a corresponding item is included ($1$) or excluded (00) from the knapsack. The total number of possible solutions is $2^{d}$, where $d$ is the number of items, reflecting the binary decision (included or excluded) for each item. More details on the experimental setup for the Knapsack problem can be found in Appendix 12.1. It is important to note that NN-GC is not readily applicable in this discrete setting because generating discrete samples with a generative model requires further modeling considerations that are beyond the scope of this paper.
Each experiment is repeated $50$ times for every method, and we calculate the average regret, along with the $5$% lower bound and $95$% upper bound, as well as the associated diversity of the suggested actions. For each trial, we sample one possible qualitative desirability function from $U$ and evaluate the regret for each policy with $m$ set to $20$. The synthetic results are summarized in Table 2.
@@
Overall, our proposed methods demonstrate significant improvements over the baseline approaches. Both NN-GC and DIS-GC yield notably lower expected regret while exhibiting reduced uncertainty in regret compared to the baselines. Our methods also achieve considerably higher diversity than QO and QO+Noise, though slightly lower than Random, underscoring the challenge of balancing diversity with optimality.
@@
Learned generative policies from the synthetic experiment with varying $\sigma$ in a $2$D space. The surface in blue indicates the PDF of learned policy, which is estimated by KDE using $10,000$ generated actions. The black dots represent $m=20$ actions suggested by the corresponding policy.
@@
Learned generative policies from the synthetic experiment with varying $m$ in a $2$D space. We set $\sigma=10$ in this comparison. The PDF of learned policy (surface in blue) is estimated by KDE using $10,000$ generated actions. The black dots represent $m$ actions suggested by the corresponding policy.
4.2 Real Results
@@
Since Atlanta’s police zones were last reconfigured in 2011, the city’s population has grown substantially, leading to increased police workload, particularly in high-growth areas such as North Atlanta (as shown in Figure 1(c)(a) and Figure 7(e)(a)). This population growth, coupled with recruitment challenges, has contributed to longer response times for high-priority 911 calls. To address these issues, we aim to reallocate $78$ geographical units within the Atlanta police system into six zones, with the goal of minimizing workload variance across the zones. The police redistricting problem can be formulated as a mixed-integer linear programming problem, where the decision variable is a $78\times 6$ binary matrix, with each entry representing the assignment of beat $i$ to zone $j$. For simplicity, the workload for each zone is calculated as the sum of workloads for all beats within that zone, and the workload of each beat is estimated using 911 calls-for-service data from 2013 to 2017. Because the real data lacks direct access to qualitative desirability, we model qualitative desirability $U$ as unknown to all models and use a GP with an exponential kernel, where $h=0.5$ and $\sigma_{U}=10^{2}$. Similar to the Knapsack problem, we use Hamming distance to measure the difference between two vectorized districting plans. The performance of all methods is assessed by measuring their best regret when $m=5$. Additionally, we extensively test our method with different variance values of $\sigma$ ranging from 00 to $10^{5}$, investigating how the choice of $\sigma$ affects regret in practice.
@@
Figure 4.2 summarizes the performance of all methods. DIS-GC achieves the lowest regret and outperforms other baseline methods when $\sigma\approx 10^{2},10^{3}$. This demonstrates that our proposed framework can effectively recommend the most desirable actions when the variance $\sigma$ and kernel function $k$ are chosen appropriately (The selection of $\sigma$ and $k$ rely on human domain expertise in practice). We also present four random plans generated by DIS-GC with $\sigma=10$, alongside the original plan used by the APD before 2019, in Figure 7(e), all of which show significant reductions in zone workload variance. In addition to minimizing workload variance and ensuring zone contiguity, the police must consider several other practical factors that are not easily captured by mathematical models when selecting a final plan. For instance, the redesigned zones should be compact and aligned with the city’s transportation infrastructure, as patrol units typically travel within a zone. A narrow or irregularly shaped zone, or one divided by highways, can increase travel time and reduce patrol efficiency and emergency response effectiveness. Notably, plan (c) closely resembles the one implemented by APD in 2019 (shown in Figure 1(c) (b)), suggesting that this method could be applicable in real-world scenarios. Further details of the experimental setup and additional numerical results are provided in Appendix 12.2.

### v2

@@ 2.2 A Tight Reformulation with Gaussian Processes
(a) Single-solution decision-making
(b) Generative curation
Comparison between classic optimal decision-making and the proposed framework. (a) In classic settings, optimization models typically focus only on observable quantitative factors with the goal of identifying a single optimal solution. (b) Generative Curation offers multiple thoughtfully curated solutions to human decision-makers, enabling them to choose the most desirable one while actively involving them in the decision-making process.
By adopting this GP-based framework, we ensure that the unknown qualitative desirability $U(a)$ is modeled in a flexible yet structured way. Specifically, utilizing Assumption 2.2, we can bound the objective in Problem 2 as followed:
@@
Problem 2 admits the following upper and lower bounds:
@@
The proof is detailed in Appendix 6. By modeling qualitative desirability with a GP, we derive upper and lower bounds that include the term $\sigma\sqrt{1-\rho[\pi]}E_{m}$. This term represents the expected gain from optimizing over $m$ solutions with uncertain qualitative desirability $U(a)$, while the gap between the upper and lower bound reflects the objective gain in using policy $\pi$ to select over $m$ solutions. In the case where $m=1$, the term $E_{m}$ becomes zero, reducing both bounds to the classical optimal policy problem: $\max_{\pi}\mathbb{E}_{\pi}[Y(A_{i})].$ This outcome confirms that when the goal is to recommend only a single solution to the decision-maker, the unobservable qualitative desirability can be effectively ignored. Figure 1(b) illustrates this contrast between classic single-solution optimization and our proposed generative curation approach.
However, when multiple solutions are to be recommended, the objective naturally evolves into a trade-off between quantitative optimality and $\sigma\sqrt{1-\rho[\pi]}E_{m}$, which can be seen as a measure of solution diversity. Specifically, $\rho[\pi]$ captures the expected correlation between the qualitative desirability components $U(A_{i})$ and $U(A_{j})$, serving as a measure of similarity or concentration among the generated solutions. When $\rho[\pi]=1$, the solutions are perfectly correlated in terms of their qualitative desirability, indicating a lack of diversity. The term $\sigma\sqrt{1-\rho[\pi]}$ in both bounds therefore incentivizes minimizing this correlation to promote diversity among the solutions.
@@
[…]
(a) Desirability
(b) $\epsilon$-optimal uniform policy
(c) $\epsilon$ vs regret
@@ 2.3 Theoretical Results
    Assume the space of actions $\mathcal{A}$ is [q,q][-q,q]. If $U(a)$ follows the Gaussian kernel $k(a,a^{\prime})=\exp\left(-\frac{(a-a^{\prime})^{2}}{2h^{2}}\right)$, let:
@@
$\text{s.t.}\;\;Y(a_{i})\geq(1-\delta)Y(a^{*}),\quad\forall a\in\text{supp}(\pi).$
@@ 2.4 Incorporating Human Preference Feedback
The framework outlined above addresses a scenario in which the algorithm provides a range of actions for the human decision-maker to choose from, particularly when there is limited information about the decision-maker’s qualitative preferences. This approach can be further extended by incorporating feedback from the decision-maker, allowing the algorithm to progressively refine its understanding of implicit preferences. By leveraging this feedback in a multi-stage process, the algorithm can adapt and improve the accuracy of future recommendations.
@@ 3.1 Deep Generative Approach
Our proposed framework allows a wide range of generative models to directly parameterize the policy $\pi$, enabling their use in solving the optimization problem presented in Problem 2. By harnessing the capabilities of cutting-edge generative models, our approach can produce a rich distribution of actions that is difficult to capture with traditional parametric models. These generative models not only optimize the objective function but also maintain a desirable level of diversity, which is critical in many real-world applications. In the experiments, we focus on a simple yet flexible generative architecture using the reparameterization trick (Kingma 2013) for clarity and ease of implementation. Specifically, instead of sampling $a$ directly from the distribution $\pi$, we express $a$ as a deterministic function of the model parameter $\theta$ and an auxiliary noise variable $\epsilon_{\text{NN}}$ drawn from a standard normal distribution. Thus, the generative process can be written as:
@@ 3.2 Iterative Curation Approach
Despite the success of generative models in various fields, applying these techniques to integer optimization problems – such as routing, assignment, and scheduling – presents unique challenges. These problems are often characterized by discrete and combinatorical structures, where the solution space is non-continuous and highly constrained. Traditional methods in integer programming focus on identifying a single optimal solution, often using techniques such as branch and bound (Land and Doig 1960) or cutting planes (Gomory 1958). These methods are not inherently designed to generate a distribution of solutions, making the direct application of the generative curation framework less straightforward.
To address this limitation, we propose an iterative approach, as outlined in Algorithm 2, that enables the generation of diverse solutions over multiple iterations. At every step, we optimize the sample estimate of the objective function in Equation (5), and over iterations, the generated solutions increasingly reflect the desired balance between diversity and optimality. As illustrated by Figure 3(c), this approach allows the model to incrementally refine the solution set. Empirically, this procedure converges to the same distribution as our optimal solution to the original problem.
@@ 4 Experiments
To ensure a fair comparison, we rerun each baseline $m$ times and select the action that gives the highest human desirability. To measure the performance of the algorithms, we adopt total regret as the key evaluation metric in our synthetic experiments, defined as:
@@
To evaluate the robustness of the methods, we also examine the $95\%$ quantile upper bound (High) and $5\%$ quantile lower bound (Low) of the regret. Furthermore, we measure empirical diversity using $\sqrt{1-\widehat{\rho}[\pi]}$, where $n$ random actions $\{a_{i}\}_{i\leq n}$ are generated from policy $\pi$, and $\widehat{\rho}[\pi]$ is computed as $1/\sigma^{2}\sum_{i=1}^{\lfloor{n/2}\rfloor}k(a_{2i},a_{2i-1})/\lfloor{n/2}\rfloor$.
@@ 4.1 Synthetic Results
Methods 11D Gaussian 22D Ackley Knapsack
$\mathbb{E}R(\pi)$ Low High $\sqrt{1-\widehat{\rho}[\pi]}$ $\mathbb{E}R(\pi)$ Low High $\sqrt{1-\widehat{\rho}[\pi]}$ $\mathbb{E}R(\pi)$ Low High $\sqrt{1-\widehat{\rho}[\pi]}$
@@
An example of the synthetic 22D desirability function in a bounded action space. The transparent surface in red represents the true desirability $\ell$ and the surface in gray represents the quantitative desirability $Y$, which is an 22D negative Ackley function. In this example, we set $\sigma=10$.
@@
    2D Ackley: We select a negative Ackley function (Ackley 2012) in a bounded 22D space $3,3]\times[-3,3]$ as the quantitative desirability $Y$ and utilize zero-mean exponential-kernel GP with $h=0.5$ for the qualitative desirability. Figure 4.1 illustrates an example of potential desirability functions with $\sigma=10$, where the quantitatively optimal action is $a^{*}=(0,0)$.
@@
    Knapsack: We apply the proposed framework to a discrete setting by solving the classic knapsack problem (Kellerer et al. 2004). This problem consists of a set of $d=10$ items, each characterized by a specific weight and value (randomly generated integers between 0 and $10$), and a knapsack with a defined capacity $20$. The objective is to determine the optimal combination of items that maximizes the total value without exceeding the knapsack’s capacity. We set $\sigma=10$ and model the qualitative desirability using a GP, where the Hamming distance is used to measure the “similarity” between two actions, and an exponential covariance kernel with $h=0.5$ is adopted. Each possible action is represented as a binary vector, where each element indicates whether a corresponding item is included ($1$) or excluded (0) from the knapsack. The total number of possible solutions is $2^{d}$, where $d$ is the number of items, reflecting the binary decision (included or excluded) for each item. More details on the experimental setup for the Knapsack problem can be found in Appendix 12.1. It is important to note that NN-GC is not readily applicable in this discrete setting because generating discrete samples with a generative model requires further modeling considerations that are beyond the scope of this paper.
Each experiment is repeated $50$ times for every method, and we calculate the average regret, along with the 55% lower bound and 9595% upper bound, as well as the associated diversity of the suggested actions. For each trial, we sample one possible qualitative desirability function from $U$ and evaluate the regret for each policy with $m$ set to $20$. The synthetic results are summarized in Table 2.
@@
Overall, our proposed methods demonstrate significant improvements over the baseline approaches. Both NN-GC and DIS-GC yield notably lower expected regret while exhibiting reduced uncertainty in regret compared to the baselines. Our methods also achieve considerably higher diversity than QO and QO+Noise, only lower to the diversity of the fully random solution Random.
@@
Learned generative policies from the synthetic experiment with varying $\sigma$ in a 22D space. The surface in blue indicates the PDF of learned policy, which is estimated by KDE using $10,000$ generated actions. The black dots represent $m=20$ actions suggested by the corresponding policy.
@@
Learned generative policies from the synthetic experiment with varying $m$ in a 22D space. We set $\sigma=10$ in this comparison. The PDF of learned policy (surface in blue) is estimated by KDE using $10,000$ generated actions. The black dots represent $m$ actions suggested by the corresponding policy.
4.2 Real-world Results
@@
Since Atlanta’s police zones were last reconfigured in 2011, the city’s population has grown substantially, leading to increased police workload, particularly in high-growth areas such as North Atlanta (as shown in Figure 6(e)(a)). This population growth, coupled with recruitment challenges, has contributed to longer response times for high-priority 911 calls. To address these issues, we aim to reallocate $78$ geographical units within the Atlanta police system into six zones, with the goal of minimizing workload variance across the zones. The police redistricting problem can be formulated as a mixed-integer linear programming problem, where the decision variable is a $78\times 6$ binary matrix, with each entry representing the assignment of beat $i$ to zone $j$. For simplicity, the workload for each zone is calculated as the sum of workloads for all beats within that zone, and the workload of each beat is estimated using 911 calls-for-service data from 2013 to 2017. Because the real data lacks direct access to qualitative desirability, we model qualitative desirability $U$ as unknown to all models and use a GP with an exponential kernel, where $h=0.5$ and $\sigma_{U}=10^{2}$. Similar to the Knapsack problem, we use Hamming distance to measure the difference between two vectorized districting plans. The performance of all methods is assessed by measuring their best regret when $m=5$. Additionally, we extensively test our method with different variance values of $\sigma$ ranging from 0 to $10^{5}$, investigating how the choice of $\sigma$ affects regret in practice.
@@
Figure 4.2 summarizes the performance of all methods. DIS-GC achieves the lowest regret and outperforms all other methods when $\sigma\in[10^{1},10^{3}]$. This demonstrates that our proposed framework can effectively recommend the most desirable actions when variance $\sigma$ and kernel function $k$ are chosen appropriately. We also present four random plans generated by DIS-GC with $\sigma=10$, alongside the original plan used by the APD before 2019, in Figure 6(e), all of which show significant reductions in zone workload variance. In addition to minimizing workload variance and ensuring zone contiguity, the police must consider several other practical factors that are not easily captured by mathematical models when selecting a final plan. For instance, the redesigned zones should be compact and aligned with the city’s transportation infrastructure, as patrol units typically travel within a zone. A narrow or irregularly shaped zone, or one divided by highways, can increase travel time and reduce patrol efficiency and emergency response effectiveness. Notably, plan (c) closely resembles the one implemented by APD in 2019 (shown in Figure 8(c)(b) in Appendix 12.2), suggesting that this method could be applicable in real-world scenarios. Further details of the experimental setup and additional numerical results are provided in Appendix 12.2.